# CMB parity asymmetry from unitary quantum gravitational physics

### v2

@@ III DQFT in de Sitter & Unitarity ($G=\hbar=c=1$)
where $r=\big{|}\frac{1}{H}\big{|}\frac{1+UV}{1-UV}$ and $\frac{V}{U}=-e^{2Ht_{s}}$. The relation between the flat FLRW Eq.4 and Eq.12 is just a simple coordinate change (Lanczos and Hoenselaers (1997); Hartman (2017)). A very practiced view of flat dS spacetime in Eq.4 is that it covers only half of the dS space. However, we must keep in mind the symmetry in Eq.6, which implies that Eq.4 can cover the entire dS space. Thus, the maximally symmetric nature of the dS spacetime sets no difference whether it is expressed in flat, closed, open FLRW coordinates or the static coordinates (Mukhanov and Winitzki (2007)). The use of mapping flat coordinates to static coordinates here is so that we can realize how unitarity can be achieved with DQFT, whereas it is lost in standard QFT, as was shown by Gibbons and Hawking in 1977 Gibbons and Hawking (1977). In Fig. 2, we show unitary quantum physics in the static dS space through the (quantum) conformal diagram, where we see the three-dimensional space divided by parity for the interior and exterior of the cosmological event horizon. We can picture here how none of the observers loses any information beyond the horizon. If any state leaves the horizon, it is only the (complimentary) direct-sum component of what is inside the horizon. This means that any observer can perfectly reconstruct the information about what is beyond the horizon. According to DQFT, every density matrix of a maximally entangled pure state is split into direct-sum of 4 components
@@
Each density matrix $\rho_{n},\,n=I,II,III,IV$ is defined by a pure state component in the corresponding (geometric) superselection sector Hilbert space $\mathcal{H}_{n}$. Since Von Neumann entropy of each component of density matrix vanishes $S_{n}=-Tr\rho_{n}\ln\rho_{n}=0$, any observer in regions I, II, III, IV (see Fig. 2) would witness pure states evolving into pure states. Thus, with DQFT we can have both unitarity and observer complementarity in dS (Kumar and Marto (2025)).
@@
Figure 2: This picture depicts the (quantum) conformal diagram of static dS spacetime Eq.12 with four regions that are related by discrete transformations and the lines with arrows indicate the arrows of time in each region. Here a quantum field operator is written as direct-sum of four components $\hat{\phi}=\frac{1}{\sqrt{2}}\left(\hat{\phi}_{I}\oplus\hat{\phi}_{II}\right)\oplus\frac{1}{\sqrt{2}}\left(\hat{\phi}_{III}\oplus\hat{\phi}_{IV}\right)$ corresponding to the regions I to IV. This is not a Penrose diagram because each point represents half of the $S_{2}$ sphere in contrast to the usual Penrose diagram. This diagram is to represent quantum physics in dS space rather than classical physics for which the usual Penrose diagram is used.
@@
According to DQFT, any state is spread across the horizon by distinct components in geometric superselection sectors. This implies that any information of the entangled state is shared by the sectorial Hilbert space, leading to pure states evolving into pure states. To be precise, the Hilbert spaces of the interior and exterior states cannot be connected by direct product ($\otimes$) but rather should be related by direct-sum ($\oplus$). This has to be like this because beyond-the-horizon time is not the same, so one cannot do the same quantum mechanics everywhere. In other words, the gravitational horizons are special, they separate different spacetimes by a boundary. Thus, they cannot be treated as a standard null surface, where quantum states can be entangled across. Furthermore, we stress that no entanglement across the horizon does not lead to firewalls, which happens in standard QFTCS but not in direct-sum QFTCS. For further analogous discussion on unitarity and no information loss in the context of the Schwarzschild black hole, see Kumar and Marto (2024). In a nutshell, DQFT offers a new understanding of dS space where any observer’s physics of the universe is complete, and no information is lost outside any observer’s universe. This is exactly what Schrödinger envisioned in 1956 for a quantum theory to achieve Erwin Schrodinger (1956). Furthermore, our achievement of DQFT echoes well with the recent classical understanding of dS through the Black Hole Universe (BHU) proposal Gaztanaga (2022); Gaztañaga (2023); Gaztanaga:2025cun.
@@ IV CMB parity asymmetry
Inflationary quantum fluctuations are the initial seeds for large-scale structure formation. It is important to understand their quantum nature to derive consistent CMB predictions. Inflation is by definition a quasi-dS expansion (Starobinsky (1980)) and this means that $\mathcal{P}\mathcal{T}$ in Eq.6 is spontaneously broken by the presence of a non-perturbative scalar field in addition to the GR’s tensor degree of freedom. Our focus here is on single-field inflation. The metric and scalar field matter quantum fluctuations are generated during inflation. The quantity which is a gauge-invariant combination of them is called the curvature perturbation ($\zeta$), and it is an effective (scalar) degree of freedom. It connects the quantum physics of inflation to CMB observations. The action for curvature perturbation at the linear order is
@@
where $H=\frac{\dot{a}}{a}$ which is nearly constant during inflation characterized by the smallness of slow-roll parameters $\epsilon=-\frac{\dot{H}}{H^{2}},\,\eta=\frac{\dot{\epsilon}}{H\epsilon}$ for $N=50-60$ number of e-foldings. The canonical variable which is a redefinition of curvature perturbation that we eventually quantize is
$v=a\zeta\dot{\phi}/H$ (15)
@@
$\hat{v}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\hat{v}_{(+)}\left(\tau,\,\textbf{x}\right)&0\\ 0&\hat{v}_{-}\left(-\tau,\,-\textbf{x}\right)\end{pmatrix}$ (16)
which generate quantum fields in a direct-sum vacuum
$|0\rangle_{\rm qdS}=|0\rangle_{\rm qdS_{I}}\oplus|0\rangle_{\rm qdS_{II}}=\begin{pmatrix}|0\rangle_{\rm qdS_{I}}\\ |0\rangle_{\rm qdS_{II}}\end{pmatrix}$ (17)
@@
in which they evolve forward and backward in time at the parity conjugate points in physical space. In analogy with Eq.6, the time reversal operation (quantum mechanically) in an expanding quasi-dS Universe is
$\tau\to-\tau\implies(t,H,\epsilon,\eta)\to(-t,-H,-\epsilon\,-\eta).$ (18)
These sign differences in the slow-roll parameters, which break the time-reversal symmetry of de Sitter spacetime, lead to corrections in the evolution of quantum fluctuations $\hat{v}_{(\pm)}$ in the parity-conjugate regions. This results in an asymmetry in the CMB temperature under spatial parity.
In spherical coordinates, the parity transformation maps a point at radial distance $r$ from angular position $(\theta,\,\varphi)$ to its antipode, $(\pi-\theta,\,\pi+\varphi)$. This transformation cannot be realized by any rotation, and thus represents a discrete global symmetry distinct from statistical isotropy. It is important to emphasize that this global spatial parity asymmetry is unrelated to the observed birefringence in the CMB polarization, which arises from parity-violating interactions with well-defined transformation properties Komatsu:2022nvu.
Figure 3: The top panels show the anti-symmetric, $\mathcal{T}^{-}(\hat{n})=\left[\mathcal{T}(\hat{n})-\mathcal{T}(-\hat{n})\right]/2$, component of the Planck 2018 CMB temperature map (Akrami et al. (2020a)) in North and South polar galactic caps. The visualization is such that the Galactic plane, shown in Fig. 4, remains obscured along the edges. Parity conjugate positions can be identified by horizontally flipping into the opposite pole, where both caps have the same information with opposing signs: $\mathcal{T}^{-}(-\hat{n})=-\mathcal{T}^{-}(\hat{n})$. The bottom panel presents the observed $\mathcal{T}=\mathcal{T}^{+}+\mathcal{T}^{-}$ maps, including both the symmetric $\mathcal{T}^{+}(\hat{n})=\left[\mathcal{T}(\hat{n})+\mathcal{T}(-\hat{n})\right]/2$ and anti-symmetric components. The evident resemblance between the top and bottom panels indicates the prevalence of the anti-symmetric component, in perfect agreement with DSI predictions Eq.27.
The CMB provides temperature fluctuations $\mathcal{T}\left(\hat{n}\right)\equiv\Delta T\left(\hat{n}\right)/T_{0}$ all over the sky $\hat{n}$:
@@
$\mathcal{T}\left(\hat{n}\right)=\sum a_{\ell m}Y_{\ell m}\left(\hat{n}\right)\quad;\quad C_{\ell}=\sum_{m}\frac{|a_{\ell m}|^{2}}{2\ell+1}$ (19)

### v4

@@ III DQFT in de Sitter & Unitarity ($G=\hbar=c=1$)
where $r=\big{|}\frac{1}{H}\big{|}\frac{1+UV}{1-UV}$ and $\frac{V}{U}=-e^{2Ht_{s}}$. The relation between the flat FLRW Eq.8 and Eq.16 is just a simple coordinate change (Lanczos and Hoenselaers (1997); Hartman (2017)). A very practiced view of flat dS spacetime in Eq. 8 is that it covers only half of the dS space. However, we must keep in mind the symmetry in Eq. 10, which implies that Eq. 8 can cover the entire dS space. Thus, the maximally symmetric nature of the dS spacetime sets no difference whether it is expressed in flat, closed, open FLRW coordinates or the static coordinates (Mukhanov and Winitzki (2007); Kumar and Marto (2025)). The use of mapping flat coordinates to static coordinates here is so that we can realize how unitarity can be achieved with DQFT, whereas it is lost in standard QFT, as was shown by Gibbons and Hawking in 1977 Gibbons and Hawking (1977). In Fig. 2, we show unitary quantum physics in the static dS space through the (quantum) conformal diagram, where we see the three-dimensional space divided by parity for the interior and exterior of the cosmological event horizon. We can picture here how none of the observers loses any information beyond the horizon. If any state leaves the horizon, it is only the (complimentary) direct-sum component of what is inside the horizon. This means that any observer can perfectly reconstruct the information about what is beyond the horizon. According to DQFT, every density matrix of a maximally entangled pure state is split into direct-sum of 4 components
@@
Each density matrix $\rho_{n},\,n=I,II,III,IV$ is defined by a pure state component in the corresponding (geometric) superselection sector Hilbert space $\mathcal{H}_{n}$. Since the Von Neumann entropy of each component of the density matrix vanishes, $S_{n}=-Tr\rho_{n}\ln\rho_{n}=0$, any observer in regions I, II, III, IV (see Fig. 2) would witness pure states evolving into pure states. Thus, with DQFT, we can have both unitarity and observer complementarity in dS (Kumar and Marto (2025)). Furthermore, this DQFT construction is also studied in the context of achieving unitarity in Rindler spacetime Kumar and Marto (2024d).
@@
Figure 2: This picture depicts the (quantum) conformal diagram of static dS spacetime Eq.16 with four regions that are related by discrete transformations and the lines with arrows indicate the arrows of time in each region. Here a quantum field operator is written as direct-sum of four components $\hat{\phi}=\frac{1}{\sqrt{2}}\left(\hat{\phi}_{I}\oplus\hat{\phi}_{II}\right)\oplus\frac{1}{\sqrt{2}}\left(\hat{\phi}_{III}\oplus\hat{\phi}_{IV}\right)$ corresponding to the regions I to IV. This is not a Penrose diagram because each point represents half of the $S_{2}$ sphere, in contrast to the usual Penrose diagram. This diagram is to represent quantum physics in dS space rather than classical physics, for which the usual Penrose diagram is used.
@@
According to DQFT, any state is spread across the horizon by distinct components in geometric superselection sectors. This implies that any information of the entangled state is shared by the sectorial Hilbert space, leading to pure states evolving into pure states. To be precise, the Hilbert spaces of the interior and exterior states cannot be connected by direct product ($\otimes$) but rather should be related by direct-sum ($\oplus$). This has to be like this because beyond-the-horizon time is not the same, so one cannot do the same quantum mechanics everywhere. In other words, the gravitational horizons are special; they separate different spacetimes by a boundary. Thus, they cannot be treated as a standard null surface, where quantum states can be entangled across. Furthermore, we stress that no entanglement across the horizon does not lead to firewalls, which happens in standard QFTCS but not in direct-sum QFTCS. For further analogous discussion on unitarity and no information loss in the context of the Schwarzschild black hole, see Kumar and Marto (2024a). In a nutshell, DQFT offers a new understanding of dS space where any observer’s physics of the universe is complete, and no information is lost outside any observer’s universe. This is exactly what Schrödinger envisioned in 1956 for a quantum theory to achieve Erwin Schrodinger (1956). The formalism of DQFT has certain similarities with ’tHooft’s recent ideas on black holes ’t Hooft (2016). However, there are some crucial differences, one of them is the construction of the geometric superselection rule that restructures a single quantum state in sectorial Hilbert spaces that represent parity conjugate regions of physical space. ’tHooft on the other hand considers anti-podal identification of time-reversal regions of spacetime with a slightly different interpretation of parity. The idea of antipodal identification was later given up by ’t Hooft because of the possibility of closed timeline curves when dynamics of spacetime are involved, so a new idea of quantum clones is proposed ’t Hooft (2022). Our framework significantly differs from this latter proposal; we speak about $\mathcal{P}\mathcal{T}$ based mirror symmetries within the description of quantum fields in the same manifold rather than creating clones in different manifolds. One can also feel some aesthetic similarities between our proposal and Boyle and Turok’s CPT symmetric universes Boyle et al. (2018, 2022) (coincidentally, it is a new understanding of the old proposal by Sakharov Sakharov (1967, 1980)). The crucial difference in our framework is that we are speaking about parity and time reversal within the same expanding universe, in contrast with Boyle and Turok’s pair of CPT-symmetric universes before and after the Big Bang. Furthermore, DQFT echoes in parallel with the recent classical understanding of dS through the Black Hole Universe (BHU) proposal Gaztanaga (2022); Gaztañaga (2023); Gaztañaga et al. (2025).
@@ IV CMB parity asymmetry
Inflationary quantum fluctuations are the initial seeds for large-scale structure formation. It is important to understand their quantum nature to derive consistent CMB predictions. Inflation is by definition a quasi-de Sitter expansion (Starobinsky (1980)) and this means that $\mathcal{P}\mathcal{T}$ in Eq. 10 is spontaneously broken by the presence of a non-perturbative scalar field in addition to the GR’s tensor degree of freedom. Our focus here is on single-field inflation. The metric and scalar field matter quantum fluctuations are generated during inflation. The quantity that is a gauge-invariant combination of them is called the curvature perturbation ($\zeta$), and it is an effective (scalar) degree of freedom ($\phi$). It connects the quantum physics of inflation to CMB observations. The perturbed action for curvature perturbation at the linear order is Gaztañaga and Kumar (2024); Baumann (2018)
@@
where $H=\frac{\dot{a}}{a}$ which is nearly constant during inflation characterized by the smallness of slow-roll parameters $\epsilon=-\frac{\dot{H}}{H^{2}},\,\eta=\frac{\dot{\epsilon}}{H\epsilon}$ for $N=50-60$ number of e-foldings. In Eq. 18, overbar indicates the background field, overdot indicates time derivative, denotes derivative with respect to conformal time, and $\partial_{i}=\frac{\partial}{\partial x^{i}}$ implies spatial gradient. The canonical variable which is a redefinition of curvature perturbation that we eventually quantize is
$v=a\zeta\dot{\bar{\phi}}/H$ (19)
@@
$\hat{v}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}\hat{v}_{+}\left(\tau,\,\textbf{x}\right)&0\\ 0&\hat{v}_{-}\left(-\tau,\,-\textbf{x}\right)\end{pmatrix}$ (20)
which can be expanded as
$\hat{v}_{\pm}=\int\frac{d^{3}k}{\left(2\pi\right)^{3/2}}\Bigg{[}c_{\left(\pm\right)\textbf{k}}{v}_{\pm,\,k}e^{\pm i\textbf{k}\cdot\textbf{x}}+c_{\left(\pm\right)\textbf{k}}^{\dagger}{v}_{\pm,\,k}^{\ast}e^{\mp i\textbf{k}\cdot\textbf{x}}\Bigg{]}$ (21)
This would also mean promoting the curvature perturbation as an operator in the following way, which involves classical rescaling
$\hat{\zeta}=\frac{H}{a\dot{\bar{\phi}}}\Bigg{|}_{\rm classical}\hat{v}$ (22)
These quantum fluctuations evolve in a direct-sum vacuum defined by
$|0\rangle_{\rm qdS}=|0_{+}\rangle_{\rm qdS}\oplus|0_{-}\rangle_{\rm qdS}=\begin{pmatrix}|0_{+}\rangle_{\rm qdS}\\ |0_{-}\rangle_{\rm qdS}\end{pmatrix},\quad c_{\pm\textbf{k}}|0_{\pm}\rangle_{\rm qdS}=0\,.$ (23)
@@
in which they evolve forward and backward in time at the parity conjugate points in physical space, governed by the mode functions $v_{\pm,\,k}$ are the solutions of
$v_{\pm,\,k}^{\prime\prime}+\left(k^{2}-\frac{{\nu}_{s}^{\left(\pm\right)2}-\frac{1}{4}}{\tau^{2}}\right)v_{\pm,\,k}^{2}=0,\quad\nu_{s}^{\pm}\approx\frac{3}{2}\pm\epsilon\pm\frac{\eta}{2}$ (24)
Recall that the choice of mode functions determines the vacuum (See Appendix C for the discussion in the context of de Sitter). The solutions of (24) are
${v}_{\pm,\,k}$ $\approx\sqrt{\frac{1}{2k}}e^{\mp ik\tau}\left(1\mp\frac{i}{k\tau}\right)\pm\left(\epsilon+\frac{\eta}{2}\right)\frac{\sqrt{\pi}}{2\sqrt{k}}\sqrt{\mp k\tau}\frac{\partial H^{(1)}_{\nu_{s}^{\pm}}\left(\mp k\tau\right)}{\partial\nu_{s}^{\pm}}\Big{|}_{\nu_{s}^{\pm}=3/2}\equiv v^{\rm dS}_{\pm,k}\left(1\pm\Delta v\right)$ (25)
where $H^{(1)}_{\nu_{s}}\left(z\right)$ is the Hankel functions of the first kind and
@@
$v_{\pm,k}^{\rm dS}=\sqrt{\frac{1}{2k}}e^{\mp ik\tau}\left(1\mp\frac{i}{k\tau}\right),\quad\Delta v=\left(\epsilon+\frac{\eta}{2}\right)\left[\frac{1}{H_{3/2}^{(1)}\left(\frac{k}{k_{\ast}}\right)}\frac{\partial H^{(1)}_{\nu_{s}}\left(\frac{k}{k_{\ast}}\right)}{\partial\nu_{s}}\Bigg{|}_{\nu_{s}=\frac{3}{2}}\right]$ (26)
The mode functions of curvature perturbation can be expressed as:
$\zeta_{\pm k}=\frac{H}{a\dot{\bar{\phi}}}v_{\pm k}=\zeta_{\pm\,k}^{\rm dS}\left(1\pm\Delta v\right)\quad;\quad\zeta_{\pm\,k}^{\rm dS}\equiv\frac{H}{a\dot{\bar{\phi}}}v_{\pm,\,\textbf{k}}^{\rm dS}$ (27)
We can see here how the extra quasi-de Sitter contribution ±$\pm\Delta v$ is anti-symmetric in parity conjugate regions of space. In analogy with Eq. 10, the time reversal operation (quantum mechanically) in an expanding quasi-dS Universe is [^2] Our notation of any quantity $Q\to-Q$ means we replace the quantities by a negative sign of that quantity.
$\tau\to-\tau\implies(t,H,\epsilon,\eta)\to(-t,-H,-\epsilon,-\eta).$ (28)
We stress that the above flip of the sign of the slow-roll parameters should not be thought of from a classical point of view. Indeed, if one does literally $t\to-t$ and $H\to-H$, the definition of slow-roll parameters does not change sign. However, what we have to keep in mind here is that quantum mechanically, the concept of time is different from the classical point of view. Time is a parameter in quantum theory and also in quantum field theory  [^3] In quantum mechanics, coordinate time is not an operator but a parameter because of the fundamental reason that time-reflection is an anti-Unitary operation Brunetti et al. (2010); Donoghue and Menezes (2019, 2021); Roberts (2022). It is worth recalling that in the usual QFT in Minkowski spacetime, which is the unification of special relativity and quantum mechanics, is built on assuming an arrow of time $t_{p}:-\infty\to\infty$, definition of positive energy ($\mathcal{E}>0)$ state $|\Psi\rangle_{t}=e^{-i\mathcal{E}t_{p}}|\Psi\rangle_{0}$ and the operators to commute for space-like distances. DQFT brings a new construction of quantum theory, which gives a new understanding of quantum physics with both arrows of time. Kumar and Marto (2025). When it comes to gravity, we encounter the dynamical nature of spacetime, so when we speak about quantum fields in curved spacetime, not only time but also any additional quantities that define the dynamics of spacetime act as parameters quantum mechanically. When we study inflationary fluctuations, which are a combination of metric and matter degrees of freedom (indeed the curvature perturbation $\zeta=-\Psi+H\frac{\delta\phi}{\dot{\bar{\phi}}}$ is a combination of Bardeen potential $\Psi$ and the inflaton field fluctuation $\delta\phi$), what we do is linearized quantum gravity (See Martin (2005) for more discussion) that is encapsulated by the following equation which also include tensor fluctuations [^4] In this paper, we only address the parity asymmetry in two-point functions of scalar fluctuations (that leads to asymmetry in the even-odd angular power spectra that is significant on large scales) according to DSI model but the DSI also predicts parity asymmetry in the tensor two point correlations as explained in Kumar and Marto (2024d).
$\delta\hat{G}_{\mu\nu}=\delta\hat{T}_{\mu\nu}$ (29)
The above sign changes (once more interpreted quantum mechanically) in the slow-roll parameters, which break the time-reversal symmetry of de Sitter spacetime, accommodate a new vacuum structure in quasi-de Sitter spacetime that lead to corrections in the evolution of quantum fluctuations $\hat{v}_{(\pm)}$ in the parity-conjugate regions. In the case of exact de Sitter spacetime, the DQFT framework leads to equal correlations for field components in the $\mathcal{P}\mathcal{T}$ conjugate regions, but in the quasi-de Sitter we do get unequal correlations because the formulation of our quasi-de Sitter vacuum Eq. 23 leads to fluctuations evolving asymmetrically at parity conjugate points in physical space. Thus, it is a pure quantum (gravitational) effect, and it has no classical analog. One can interpret this result as inflation breaks spontaneously the $\mathcal{P}\mathcal{T}$ symmetry of de Sitter space, implying the quantum fluctuations to evolve asymmetrically at parity conjugate points (see Fig. 3 in Gaztañaga and Kumar (2024) or Fig.1 above). As we will show, this results in an asymmetry in the CMB temperature under spatial parity. Following the two (scalar) power spectra defined with respect to the two quasi-de Sitter vacua Eq. 23 that correspond to parity conjugate regions of physical space. From Eq.27 we have:
$\mathcal{P}_{\zeta\pm}=\int d^{3}re^{-i\textbf{k}.\textbf{r}}{}_{\rm qdS}\langle 0_{\pm}|\hat{\zeta}_{\pm}\left(\pm\tau,\,\pm\textbf{x}\right)\hat{\zeta}_{\pm}\left(\pm\tau,\,\pm\textbf{x}^{\prime}\right)|0_{\pm}\rangle_{\rm qdS}\approx P_{\zeta}\left(1\pm\Delta\mathcal{P}_{v}\right)$ (30)
where $r=|\textbf{x}-\textbf{x}^{\prime}|$ and $\mathcal{P}_{\zeta}$ is the near scale invariant part of the power spectrum $P_{\zeta}$:
$P_{\zeta}=\frac{H^{2}}{8\pi\epsilon}\approx A_{s}\left(\frac{k}{k_{\ast}}\right)^{n_{s}-1}\,.$ (31)
where $n_{s}=1-2\epsilon-\eta$. In standard inflation (SI) calculations, only the term $+$ (i.e., the wave with $e^{-ik\tau}$ in Eq.26) is considered, and the slow-roll corrections are usually neglected. This resulting in the (near) scale-invariant power spectrum (See Appendix G in Gaztañaga and Kumar (2024) for more details). In DSI, the correction, $\Delta\mathcal{P}_{v}$ comes with positive and negative contributions to the (near) scalar-invariant part Eq. 31 at parity conjugate points because we apply (quantum mechanically) the time reversality Eq. 28 to the quasi-de Sitter vacuums $|0_{\pm}\rangle_{\rm qdS}$ Eq. 23 at the parity conjugate regions.
From the latest Planck data Akrami et al. (2020a) $A_{s}=2.2\times 10^{-9}$ and the value of spectral index $n_{s}$ corresponding to the pivot scale $k_{\ast}=a_{\ast}H_{\ast}=0.05{\rm Mpc}^{-1}$ is
$n_{s}=1-2\epsilon_{\ast}-\eta_{\ast}\approx 0.9634\pm 0.0048\quad\left(\rm{PlanckTT+TE+EE}\right)$ (32)
In DQFT inflation (DSI), we include both terms as a direct-sum, which results in Eq.27 and leads to Eq.30 with:
$\Delta\mathcal{P}_{v}=\left(1-n_{s}\right)\operatorname{Re}\left[\frac{2}{H_{3/2}^{(1)}\left(\frac{k}{k_{\ast}}\right)}\frac{\partial H^{(1)}_{\nu_{s}}\left(\frac{k}{k_{\ast}}\right)}{\partial\nu_{s}}\Bigg{|}_{\nu_{s}=\frac{3}{2}}\right].$ (33)
Note that the $\pm$ in Eq. 30 comes from the difference in the quantum vacuum Eq. 23 in the parity conjugate regions emerging from the time reversal operation Eq. 28. In other words, the additional term $\Delta\mathcal{P}_{v}$ in Eq. 30 corresponds to a purely anti-symmetric contribution that changes sign at parity conjugate points. In spherical coordinates, the parity transformation maps a point at radial distance $r$ from angular position $(\theta,\,\varphi)$ to its antipode, $(\pi-\theta,\,\pi+\varphi)$. This transformation cannot be realized by any rotation, and thus represents a discrete global symmetry distinct from statistical isotropy. It is important to emphasize that this global spatial parity asymmetry is unrelated to the observed birefringence in the CMB polarization, which arises from parity-violating interactions with well-defined transformation properties Komatsu (2022).
Figure 3: The top panels show the anti-symmetric, $\mathcal{T}^{-}(\hat{n})=\left[\text{T}(\hat{n})-\text{T}(-\hat{n})\right]/2$, component of the Planck 2018 CMB temperature map (Akrami et al. (2020b)) in North and South polar galactic caps. The visualization is such that the Galactic plane, shown as a grey mask in Fig. 5, remains obscured along the edges. Parity conjugate positions can be identified by horizontally flipping into the opposite pole (as indicated by the dashed arrow example) where we find similar structures with opposing signs: $\mathcal{T}^{-}(-\hat{n})=-\mathcal{T}^{-}(\hat{n})$. The bottom panel presents the observed $\mathcal{T}=\mathcal{T}^{+}+\mathcal{T}^{-}$ maps, including both the symmetric $\mathcal{T}^{+}(\hat{n})=\left[\text{T}(\hat{n})+\text{T}(-\hat{n})\right]/2$ and anti-symmetric components. The evident resemblance between the top and bottom panels indicates the prevalence of the anti-symmetric component, in perfect agreement with DSI predictions Eq.38.
IV.1 CMB temperature sky fluctuations
On very large, super-horizon scales, the dominant contribution to CMB anisotropies arises from the Sachs-Wolfe effect Sachs and Wolfe (1967), which corresponds to the gravitational redshift of the CMB photons due to small variations in curvature perturbations. This results in a one to one correspondence between temperature and metric fluctuations: $\Delta T\left(\hat{n}\right)\sim\zeta$. The observed CMB temperature fluctuation map $\text{T}\left(\hat{n}\right)\equiv\Delta T\left(\hat{n}\right)/T_{0}$ over the sky position $\hat{n}$ can be decomposed into spherical harmonics:
$\text{T}\left(\hat{n}\right)=\sum a_{\ell m}Y_{\ell m}\left(\hat{n}\right),\quad a_{\ell m}=\int\sin\theta\ d\theta\ d\varphi\ \text{T}(\hat{n})\ Y_{\ell m}^{\ast}\left(\theta,\,\varphi\right);\quad C_{\ell}=\sum_{m}\frac{|a_{\ell m}|^{2}}{2\ell+1}$ (34)
For a Gaussian field, the statistical properties of the Fourier-like harmonic coefficients $a_{\ell m}$ are characterized by the angular power spectrum $C_{\ell}$, as defined above. For antipodal (parity-conjugate) sky positions $(-\hat{n})$, the spherical harmonics $Y_{\ell m}$ satisfy:
$Y_{\ell m}(-\hat{n})=Y_{\ell m}\left(\pi-\theta,\,\pi+\varphi\right)=\left(-1\right)^{\ell}Y_{\ell m}\left(\theta,\,\varphi\right)=\left(-1\right)^{\ell}Y_{\ell m}(\hat{n}).$ (35)
The additional DSI term ±$\pm\Delta v$ in Eq. 27 corresponds to a purely antisymmetric contribution that changes sign at parity-conjugate points (due to time-reversal operation). This translates into an additional purely antisymmetric temperature fluctuation $\delta T$ in the DSI model superimposed to the standard inflation (SI) contribution $\text{T}^{SI}(\hat{n})$:
$\text{T}^{DSI}\left(\hat{n}\right)=\text{T}^{SI}(\hat{n})\left(1+\delta\text{T}\right)\quad;\quad\delta\text{T}\left(\hat{n}\right)=-\delta\text{T}\left(-\hat{n}\right)\,\implies\delta\text{T}\left(\hat{n}\right)=\sum_{\ell,\,m}\left(-1\right)^{\ell+1}\delta a_{\ell m}Y_{\ell m}\left(\theta,\,\varphi\right)$ (36)
This shows that when the fluctuations $\delta a_{\ell m}$ are smoothed and do not vary significantly from one multipole to the next, the additional antisymmetric DSI contribution results in an alternating negative (for even multipoles) and positive (for odd multipoles) modulation on top of the SI signal.
In the Sachs-Wolfe regime, we can relate the 2D angular power spectrum $C_{\ell}$ to the metric 3D metric power spectrum $P_{\zeta}(k)$ by a simple projection (see e.g. Durrer (2020)):
$C_{\ell}=\frac{2}{9\pi}\int\frac{dk}{k}\mathcal{P}_{\zeta}\left(k\right)j_{\ell}^{2}\left(k/k_{s}\right)\,$ (37)